28.
Livestock feed. Soybean meal is 16% protein and corn meal is 9% protein. How many pounds of each should be mixed to get a 350-lb mixture that is 12% protein?

Hope someone can help me with this, thanks.

Answer by nerdybill(7384)

Livestock feed. Soybean meal is 16% protein and corn meal is 9% protein. How many pounds of each should be mixed to get a 350-lb mixture that is 12% protein?
.
Let x = amount (lbs) of 16%
then
350-x = amount (lbs) of 9%
.
.16x + .09(350-x) = .12(350)
.16x + 31.5-.09x = 42
.07x + 31.5 = 42
.07x = 10.5
x = 150 lbs (of 16% protein)
.
9% protein:
350-x = 350-150 = 200 lbs